UT 20150129 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, WSeifert, APramskiy, DThompson (LBTO)
Observer AO-Team: SEsposito, GAgapito, EPinna, APuglisi
Telescope Operator: GBechetti
LBTO Support: DMiller

Plan for the night

On sky:

- Do first closed-loop observations with LUCI2+FLAO on sky
- Start with brightest sources first, close to ROTangle = 205deg
- If successfull, do some offset tests in HeI, FeII and Br_gam
- Evaluate also PSF-variations across FOV

If closed-dome:

- Flexure compensation verification N30 camera

Summary

- First AO-images on sky: Strehl of 45% (HeI), 75% (FeII) and 75% (Br_gam) achieved

Preparations

21:40 (28.1.) - MCE disconnect: 4 restarts all while instrument units have NOT been moved
=> exchange with LUCI1 spare

Details

All times are UT
Files are luci2.20150129.NNNN.fits

00:00 ARGOS calibration Unit put in for some NCPA work

00:46 Sunset, partly cloudy but increasing coverage, T_ambient = 3.7, T_mirror = 3.5, Hum = 10%

01:15 Pretty much cloudy outside, we nevertheless give it a try, Gino opens

01:27 Check pointing and collimate at WT10_342, takes a little bit (1 hour)...

01:54 Star lost for whatever reason, check pointing, done, and collimated

Yes we are on sky!!!

02:30 Point to Hip15925 (R_mag = 6.4), PA=-10 for ROTangle = 200, DIMM seeing 0.9"

02:35 Sent adaptive preset, AO-star is at 975,1054 (N3.75camera), N30: 430,880

__First images with N30 on sky

02:44 Subwin readout, FeII, DIT=0.17sec, ROTangle = 202deg, #0001 no AO, #0002 loop closed (wo NCPA), Strehl 45% , #0003 with NCPA Strehl 62%
Rotator stopped @ 205deg, NDIT = 50 = 8.5sec Strehl 66% ,
without mirror ventilation Strehl 75%!!!

LUCI2_N30_FeII.jpg

02:48 Do offset, doesn't work properly, yes, in the end

03:40 Switch to Br_gam #0005-54 Strehl 75%!!! , seeing 0.75-0.9"

LUCI2_N30_Brgam.jpg

03:59 Switch to HeI #0055+ Strehl 35%, up to 45%

04:05 Move back to FeII, do some tests, + SL-image

Check of NCPA correction vs. ROTangle => interpolation works convincingly

04:10 Preset to HIP 30762, R_mag = 6.4 @ PA=50

05:00 Clouds come in and becoming increasingly thicker frown, sad smile

Star lost

Time for the Arcetri group to have their diner smile

05:30 100% cloud coverage

07:20 A gap in the clouds, back on sky, Gino checks pointing

07:25 Preset to HIP 41798, R_mag = 6.1, PA=0, FeII, collimate, than AO preset, failed

07:35 Changed to PA = -30, stopped rotator

07:50 Clouds passing, stopped

08:20 Sky a bit better, test offsets,

09:00 Overcast, telescope parked

10:00 ARGOS calibration unit put in, check orientation of WFS LUCI2 wrt to FLAO

12:00 Still overcast, night over

-- %USERSIG{JochenHeidt - 2015-1-29}%
I Attachment Action Size Date Who Comment
LUCI2_N30_Brgam.jpgjpg LUCI2_N30_Brgam.jpg manage 243 K 30 Jan 2015 - 08:27 JochenHeidt LUCI2+FLAO Br gamma, Strehl 75%
LUCI2_N30_FeII.jpgjpg LUCI2_N30_FeII.jpg manage 221 K 30 Jan 2015 - 08:21 JochenHeidt LUCI2+FLAO FeII, Strehl 75%
Topic revision: r10 - 19 Oct 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback